OpenStudy (anonymous):

Find the equation of the line that passes through (4, 5) with a slope of 2.

OpenStudy (texaschic101):

There is a couple of ways to do this... either by using the point-slope formula, or you can also use y = mx + b form. Which one do you want to use ?

OpenStudy (anonymous):

y=mx+b

OpenStudy (texaschic101):

ok (4,5) x = 4 and y = 5 slope(m) = 2 now we sub y = mx + b 5 = 2(4) + b 5 = 8 + b 5 - 8 = b -3 = b your equation is : y = 2x - 3

OpenStudy (texaschic101):

so basically, when you plug in your stuff, you are looking for b, which is your y intercept. If you were to use point-slope formula : y - y1 = m(x - x1) and solved for y, you would get the same answer